OpenStudy (anonymous):

27x + 9 = 13x + 51 Solve for x

alones (alones):

27x+9=13x+51 Step 1: Subtract 13x from both sides. 27x+9−13x=13x+51−13x 14x+9=51 Step 2: Subtract 9 from both sides. 14x+9−9=51−9 14x=42 Step 3: Divide both sides by 14. 14x/14=42/14 x=3